OpenStudy (anonymous):

Alicia estimated that the number of horses living in a stable was 26. The actual number of horses living in the stable was 30. What is the percent error in Alicia’s estimate? If necessary, round to the nearest percent. ? % @muscrat123 can you help (Medal)

OpenStudy (inexilewetrust):

I am not exactly sure how to do this. I have an idea but I am not fully sure. Do you know how to do find percentages?

OpenStudy (anonymous):

ya kinda

OpenStudy (inexilewetrust):

You divide the numerator (top) by the denominator (bottom) Your fraction in this case would be 26/30. What would this give you?

OpenStudy (anonymous):

0.8666 repeat

OpenStudy (anonymous):

is that my answer

OpenStudy (calculusxy):

\[\frac{ greater - lesser }{ actual } \times 100\]

OpenStudy (calculusxy):

So your greater value is 30 and the lesser value is 26. But the actual number was 30. Here is the new expression. \[\frac{ 30 - 26 }{ 30 } \times 100\]

OpenStudy (calculusxy):

And we multiply it by 100 to get the percentage from the decimal.

OpenStudy (calculusxy):

Hope this helps you :)

OpenStudy (anonymous):

13.333333333333333

OpenStudy (anonymous):

right @InExileWeTrust

OpenStudy (inexilewetrust):

Sweet. And sorry, had something else to care of. Do you get it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you